OpenStudy (juan337492):

The area of a rectangle is 440 square centimeters. If the perimeter is 84 centimeters, find the length and width of the rectangle. (w)(84/2-w)=440.

OpenStudy (danjs):

You can make two equations with Length L and width W. Area = Length x Width Perimeter = 2*Length + 2*Width

OpenStudy (danjs):

440 = L * W and 84 = 2*L + 2*W

OpenStudy (danjs):

Can you solve those for L and W?

OpenStudy (juan337492):

L=20 w=22 okay thanks